a short magnet of magnetic moment 2Am square is lying in a horizontal plane with its north pole pointing 60 degree east of north.find the net horizontal magnetic field at a point north of the magnet 0.2m away from it.horizontal component of earths magnetic field is 0.3 G.

can u pls answer it?

Grade:12

1 Answers

ROSHAN MUJEEB
3 years ago
Magnetic moment =2 Am2, the value of the earth magnetic field is wrong it should be power -4
netfield=0.3*10−4T+10−7*20.23cos60=0.3*10−4T+0.125*10−4=0.425*10−4T